OpenStudy (anonymous):

solve -2.7 * (-8.3) i have gotten so far as to coming up with either 6.5 or 5.6 i am literally stuck

OpenStudy (anonymous):

do you just want answer?

OpenStudy (anonymous):

yes i cannot figure this one out

OpenStudy (anonymous):

:) please

OpenStudy (anonymous):

83/35 i do not know the work though

OpenStudy (anonymous):

oh just do -8.3 * -2, then divide that answer by 7 and you get 83/35

OpenStudy (anonymous):

ok that is where i am stuck it is either 6.5 or 5.6

OpenStudy (anonymous):

= 16.6/7 = 83/35

hero (hero):